# Hierarchical Reinforcement Learning with Augmented Step-Level Transitions for LLM Agents

### 07 Apr 2026 v1

@@ Abstract
Shuai Zhen1, Yanhua Yu1*, Roupei Guo2, Nan Cheng2, Yang Deng3 1Beijing University of Posts and Telecommunications 2China Mobile Group Design Institute Co., Ltd 3Singapore Management University *Correspondence: yuyanhua@bupt.edu.cn
@@ 1 Introduction
Large language models (LLMs) have demonstrated remarkable capabilities as autonomous agents in sequential decision-making tasks, exhibiting sophisticated reasoning and planning abilities across diverse interactive environments (Wang et al., 2023; Yao et al., 2022; Li et al., 2022). To further enhance the effectiveness of autonomous agents, reinforcement learning (RL) offers a principled mechanism for enhancing agent decision-making capabilities  (Xu et al., 2024; Pang et al., 2024; Peiyuan et al., 2024). Unlike supervised approaches that rely solely on fixed demonstrations (Zeng et al., 2024a; Lin et al., 2023), RL enables agents to refine policy through environmental interaction and reward feedback, thereby discovering more effective strategies that generalize beyond training distributions.
Despite this progress, most RL-based LLM agents adopt a history-conditioned formulation, where policies are conditioned on increasingly long sequences of past observations and actions rather than a compact representation of the current decision state. This design choice is largely inherited from sequence-modeling perspectives: LLM agents are built on Transformer architectures (Vaswani et al., 2017), and recent RL formulations cast decision-making as trajectory or sequence prediction (Chen et al., 2021; Janner et al., 2021; Ni et al., 2023). While long histories can help infer latent states in partially observable environments, conflating long-horizon decision-making with long-context conditioning introduces fundamental limitations. Attention-based inference scales quadratically with context length, and unfiltered histories accumulate redundant or irrelevant information that can obscure decision-critical signals and degrade reasoning quality (Zhou et al., 2025; Cherepanov et al., 2023). Importantly, long-context conditioning is a modeling choice rather than a necessity of reinforcement learning.
@@ 4.2 Results
Table 1: Main Results. Performance comparison across three backbone models on ScienceWorld and AlfWorld benchmarks. indicates prompt-based methods without model parameter update, while represents fine-tuning approaches using LoRA. $\uparrow$ denotes the performance improvement of STEP-HRL compared to the best results among the baselines.
@@ 4.4 Analysis on Efficiency
As shown in Figure 3, we analyze the inference-time efficiency of STEP-HRL in comparison with standard RL and HRL on an ALFWorld task. For a fair comparison, all methods are evaluated under the same observation and action sequence. The task is decomposed into four subtasks and requires a total of 29 environment steps to complete.
@@ 4.5 Analysis on Offline RL
Figure 4 presents a sensitivity analysis of the proposed offline RL procedure with respect to key algorithmic and data-related factors. We observe that the advantage temperature $\beta$ plays a critical role in balancing update aggressiveness and stability: among the tested values, $\beta=0.95$ consistently yields the highest final performance, while both smaller and larger temperatures lead to inferior convergence. Similarly, varying the expectile parameter $\tau$ reveals that a higher expectile (e.g., $\tau=0.9$) provides more effective value estimation and results in stronger policy improvement compared to lower settings.
@@ Appendix B Training Details
advantage weighted factor $\beta$ 0.95
@@ Appendix C Evaluation Details
We further report model performance on each individual task family in ScienceWorld and ALFWorld. Since the result distributions are similar across different backbone models, we only present results for Llama-3-8B in Tables 5 and 6.

### 15 Apr 2026 v2

@@ Abstract
Shuai Zhen1, Yanhua Yu1*, Roupei Guo2*, Nan Cheng2, Yang Deng3 1Beijing University of Posts and Telecommunications 2China Mobile Group Design Institute Co., Ltd 3Singapore Management University *Correspondence: yuyanhua@bupt.edu.cn, guoruopei@cmdi.chinamobile.com
@@ 1 Introduction
Large language models (LLMs) have demonstrated remarkable capabilities as autonomous agents in sequential decision-making tasks, exhibiting sophisticated reasoning and planning abilities across diverse interactive environments (Wang et al., 2023; Yao et al., 2022; Li et al., 2022; Ren et al., 2025). To further enhance the effectiveness of autonomous agents, reinforcement learning (RL) offers a principled mechanism for enhancing agent decision-making capabilities  (Xu et al., 2024; Pang et al., 2024; Peiyuan et al., 2024). Unlike supervised approaches that rely solely on fixed demonstrations (Zeng et al., 2024a; Lin et al., 2023), RL enables agents to refine policy through environmental interaction and reward feedback, thereby discovering more effective strategies that generalize beyond training distributions.
Despite this progress, most LLM agents adopt a history-conditioned formulation, where policies are conditioned on increasingly long sequences of past observations and actions. This design choice is largely inherited from sequence-modeling perspectives: LLM agents are built on Transformer architectures (Vaswani et al., 2017), and recent RL formulations cast decision-making as trajectory or sequence prediction (Chen et al., 2021; Janner et al., 2021; Ni et al., 2023). While long histories can help infer latent states in partially observable environments, conflating long-horizon decision-making with long-context conditioning introduces fundamental limitations. Attention-based inference scales quadratically with context length, and unfiltered histories accumulate redundant or irrelevant information that can obscure decision-critical signals and degrade reasoning quality (Zhou et al., 2025; Cherepanov et al., 2023). Importantly, long-context conditioning is a modeling choice rather than a necessity of reinforcement learning.
@@ 4.2 Results
Table 1: Main Results. We report performance comparisons for three backbone models on the ScienceWorld and ALFWorld benchmarks. corresponds to prompt-based methods without updating model parameters, and corresponds to LoRA-based fine-tuning approaches. $\uparrow$ indicates the relative improvement of STEP-HRL compared to the best results among the baselines.
@@ 4.4 Analysis on Efficiency
Token Usage Efficiency.
As shown in Figure 3, we analyze the token efficiency of STEP-HRL in comparison with standard RL (i.e., a normal ReAct-style agent) and HRL on an ALFWorld task. For a fair comparison, all methods are evaluated under the same observation and action sequence. The task is decomposed into four subtasks and requires a total of 29 environment steps to complete.
@@
Wall-Clock Inference Efficiency.
We further conduct analysis on wall-clock inference latency. As shown in Table 3, STEP-HRL is slightly slower than normal agent framework in early steps. This is because STEP-HRL contains three policies and therefore performs more forward passes per step. Under KV-cache decoding, this introduces additional prefills, and local-progress generation typically produces more tokens than primitive actions.
However, as the interaction history grows, the latency of the history-conditioned baseline increases, whereas STEP-HRL remains stable. This suggests that STEP-HRL is less sensitive to history length and scales more favorably in long-horizon settings.
Table 3: Per-step inference latency on a single A100 GPU in the ALFWorld Pick&Place. $\star$ denotes subtask generation, which introduces an additional forward pass.
Step Ours (s) Normal (s)
1 1.8 $\pm$ 0.3 1.3 $\pm$ 0.3
10 1.6 $\pm$ 0.2 1.4 $\pm$ 0.2
20 1.5 $\pm$ 0.2 1.6 $\pm$ 0.2
30 1.5 $\pm$ 0.3 1.7 $\pm$ 0.3
40 1.6 $\pm$ 0.2 1.9 $\pm$ 0.2
@@ 4.5 Analysis on Offline RL
Figure 4 presents a sensitivity analysis of the proposed offline RL procedure with respect to key factors. We observe that the advantage temperature $\beta$ plays a critical role in balancing update aggressiveness and stability: among the tested values, $\beta=0.95$ consistently yields the highest final performance, while both smaller and larger temperatures lead to inferior convergence. Similarly, varying the expectile parameter $\tau$ reveals that a higher expectile (e.g., $\tau=0.9$) provides more effective value estimation and results in stronger policy improvement compared to lower settings.
@@ 6 Conclusion
Acknowledgments
The work was supported by the National Natural Science Foundation of China (Grant No. U22B2019), China Scholarship Council (No. 202406470057) and Beijing University of Posts and Telecommunications-China Mobile Communications Group Co., Ltd. Joint Institute.
@@ Appendix B Training Details
advantage weighted factor $\beta$ 1
@@ Appendix C Evaluation Details
We further report model performance on each individual task family. Since the result distributions are similar across different backbone models, we only present results for Llama-3-8B in Tables 6 and 7.